why we have to multiplt by 5/18 only to convert km/h into m/s /
 

Grade:11

1 Answers

Soumya Ranjan Mohanty
6 years ago
It is because 5/18 = 1000/3600.We do that as, 1 km= 1000 m and 1 hr = 3600 seconds.. So to convert km/hr to m/s we use 5/18.We can instead use 1000/3600 also.